OpenStudy (anonymous):

Molly's basic cell phone plan charges $35.99 for service plus $0.20 per minute beyond the included 250 minutes in the plan. How much will it cost Molly if she used her cell phone 440 minutes in one month?

OpenStudy (anonymous):

35.99 for the original fee. 38.00 for the overages. 73.99 total.

OpenStudy (anonymous):

440 - 250 = 190 190*0.20 = 38.00 38.00 +35.99 = 73.99